OpenStudy (anonymous):

Find all radian solutions: 4sin^4 + 4cosx - 5 = 0

OpenStudy (anonymous):

4sin^4x + 4cosx - 5 = 0 ?

OpenStudy (anonymous):

How do you solve to find all radian solutions for that equation?

OpenStudy (tkhunny):

Are you sure the middle term is \(\cos(x)\) and not \(\cos^{2}(x)\)?

OpenStudy (anonymous):

Yes, i'm sure it's 4cosx. I'm not quite sure how to solve for the radian this way.

OpenStudy (tkhunny):

But you're not sure that the exponent on the first term is 4. How about 2?

OpenStudy (anonymous):

The exponent on the first term is 2 not sure. I made a mistake.

OpenStudy (tkhunny):

Why don't we move the discussion to the othe rthread where this error has not been made?

OpenStudy (anonymous):

alright, sure!